OpenStudy (katielong):

f(x)=3x-5 work out f(3)-f(1) This was a question in the exam i had today. just curious about it, thanks =)

OpenStudy (anonymous):

f(3) = 3*3-5 = 4 f(1) = 3*1 -5 = -2 f(3)-f(1)= 4-(-2) =6

OpenStudy (katielong):

wooo! thats what i got! thanks =)

OpenStudy (anonymous):

f(x)=3x-5 Substitute the value of x i.e 3 and 1 in both the equations f(3)= 3*3-5 = 4 f(1) = 3*1 -5 = -2 f(3)-f(1)= 4-(-2) =6